OpenStudy (josedavid):

Write the equation of the line that passes through (3, 4) and (2, –1) in slope-intercept form. y = 3x – 7 y = 3x – 5 y = 5x – 11 y = 5x – 9 please help

OpenStudy (josedavid):

@Abbles @agent0smith @AloneS @ShadowLegendX @sweetburger

OpenStudy (josedavid):

@Bluff @NatalieMagadan1998 @andrejrusa123 @legomyego180

OpenStudy (abbles):

Do you know how to find the slope? Hint: y2-y1/x2-x1

OpenStudy (josedavid):

nope i think is a

OpenStudy (abbles):

Well, for the slope you would get: 4-(-1)/3-2 5/1 The slope is 5. Slope intercept form is y = mx + b Where m is the slope and b is the y-intercept.

OpenStudy (abbles):

So if you know the slope is 5, what do you think the answer is?

OpenStudy (josedavid):

c

OpenStudy (abbles):

Why do you say that?

OpenStudy (josedavid):

becuase ot has 5

OpenStudy (abbles):

C and D both have a 5, yeah?

OpenStudy (abbles):

You're on the right track. It will be either C or D, you know that A and B are out because the slope for those two equations is 3, not 5.

OpenStudy (abbles):

Next you need to find b, which is the y-intercept. Do that by plugging in a set of coordinates - we'll use (3,4) and solving for b. y = mx + b 4 = 5(3) + b Do you see what I did there? I plugged the x-coodinate in for x, the y coordinate in for y, and the 5 for the slope. Now we need to solve for b to find the y-intercept. Do you think you can do that?

OpenStudy (josedavid):

so c

OpenStudy (abbles):

Did you solve for the y-intercept?

OpenStudy (abbles):

Solve this for b and tell me what you get. 4 = 5(3) + b

OpenStudy (josedavid):

11

OpenStudy (josedavid):

i go it correct thank u and it was c

OpenStudy (abbles):

Your welcome